# references.bib A Learning-based Quadcopter Controller with Extreme Adaptation

### v1

@@ I Introduction
The agile nature of quadcopters and the necessity for precise control in dynamic environments create a unique context for exploring control strategies. Model-based controllers for quadcopters generally rely on estimates of the vehicle’s properties, including inertia, motor constants, and other parameters. Notable examples include sliding mode control [ZHENG2014slidingmode] and PID controllers [mellinger2012trajectory]. Once these parameters are estimated, the controller typically requires iterative tuning through successive experiments to refine its performance. However, inaccuracies in parameter estimation can directly lead to execution errors in controller commands. Furthermore, any modification to the vehicle, such as attaching an extra payload, could lead to suboptimal performance without repeating the estimation and tuning processes. Such significant engineering effort could be eased with a universal controller that does not require specialized tuning.
In this work, we propose a learning-based low-level controller designed to control a variety of quadcopters with notable differences in mass, size, propellers, and motors. Our controller is also capable of rapidly adapting to unknown in-flight disturbances such as off-center payloads, motor failures, and wind.
@@ I-A Related Work
The development of fixed-parameter controllers, while foundational, is inherently limited by their lack of real-time adaptivity to model uncertainties and disturbances. Adaptive control techniques were introduced to address these unpredicted variations in a system. One of the initial contributions in this field was the model reference adaptive controller (MRAC), an extension of the well-known MIT-rule [MAREELSMit]. The empirical success of MRAC led to the development of control [cao2008design, hovakimyan2010l1], which offers a promising solution by estimating the differences between the nominal state transitions predicted by the reference model and those observed in practice. Such differences are compensated by allocating a control authority proportional to the disturbance, effectively driving the system back to its reference behavior.
@@
However, the performance of the classic formulation degrades when the observed transitions deviate greatly from the (usually linear) reference model. This limitation is critical in scenarios involving large variations between different quadcopters’ dynamics, as the underlying assumptions on locally linear disturbances are generally not fulfilled. Recent work has attempted to overcome these limitations by combining adaptive control with nonlinear online optimization schemes, such as model predictive control (MPC) [hanover2021performance, pravitra2020, pereida2018adaptive]. These methods have achieved impressive results, but still require explicit and accurate knowledge of the reference model, which is crucial for adaptation.
@@
Recent advances in data-driven control strategies have shown promising results for quadcopter stabilization [hwangbo2017control, koch2019reinforcement], or waypoint tracking flight [song2021autonomous, kaufmann2022benchmark], as well as agile racing against human pilots [kaufman2023racingdrone]. Model-free reinforcement learning in [kaufman2023racingdrone] has demonstrated impressive adaptability to unmodeled disturbances, such as blade flapping effects. Combining data-driven methods with model-based control designs has also been proposed to leverage the guaranteed adaptivity and robustness offered by model-based control. For instance, some studies have learned policies from model-based methods like MPC through imitation learning [tagliabue2021demonstrationefficient, zhao2023efficient]. Another approach is augmenting the learned controller with classical adaptive control designs during deployment for fast disturbance estimation and online adaptation [neuralfly, huang2023datt]. Despite these advancements, these methods remain tailored to specific platforms. Transferring the same controller to another vehicle typically requires retraining or fine-tuning the policy, along with data collection for the new vehicle.
@@
Zero-shot adaptation across different vehicles has been demonstrated on quadrupeds [feng2022genloco], highlighting the versatility of learning-based methods. However, this generalized control relies on existing internal motor control loops rather than directly adapting at the motor level. In the case of quadcopters, the variation in actuators between different vehicles is particularly significant, with motor constants potentially differing by orders of magnitude. Consequently, effective adaptation across different quadcopter platforms must address motor-level differences directly. Additionally, the high-frequency nature of motor control increases the risk of crashes due to inadequate adaptation. These factors, the substantial variation in actuators and the high-frequency nature of motor control, pose significant challenges to applying previous methods of adaptive trajectory control to the problem of extreme adaptation across quadcopters.
@@ I-B Our Contributions
In this work, we present a general framework for learning low-level adaptive controllers that are effective across a wide range of quadcopters. Similar to prior research in learning-based control for aerial robots [kaufmann2020RSS, kaufmann2022benchmark, eschmann2024learning, molchanov2019sim, hwangbo2017control, becker2020learning], we train the policy entirely in simulation using reinforcement learning and deploy it directly to the real world without fine-tuning (i.e., zero-shot deployment).
While previous works typically rely on slight parameter randomizations ( 20%) around a nominal model, our approach must handle parameter variations thousands of times larger. This presents a significant challenge for reinforcement learning, as such a broad range of variations can hinder optimization convergence.
To address this challenge, we build on our earlier conference work on learning-based low-level control [zhang2023universal] and introduce three key technical innovations: (1) a dual training strategy that combines behavior cloning from specialized model-based controllers and model-free reinforcement learning. This combination effectively handles the challenges of training a low-level controller due to its high-frequency nature and the low informational density of observations. (2) A specifically designed reward to provide the low-level controller with direct feedback for quick adjustments, allowing it to perform more agile maneuvers, and (3) a designed-informed reward and domain randomization method to ensure that the variations in quadcopter designs during training are consistent with real-world constraints. These innovations eliminate constraints on slow flight and accurate state estimation, and widens the range of vehicles our policy can fly. Our approach significantly outperforms existing baselines. In addition, it enables the controller to adapt to out-of-distribution quadcopters up to 16x wider than the training set and to disturbances for which it was not explicitly trained, such as wind.
@@ II Method
We present our methodology for learning an adaptive low-level controller for various quadcopters. For clarity, we provide a list of symbols and notation used throughout the paper in Table I.
Notation | |||
@@
Commanded quantity | Estimated quantity | ||
@@
Individual motor speeds | Individual motor PWM commands | ||
@@ II-A Control Structure
Cascade control systems are instrumental in managing complex dynamic systems by decomposing them into a hierarchy of simpler nested subsystems. In this structure, the high-level component focuses on high-level tasks such as trajectory planning, while the low-level component acts as the inner control loop to execute the commands from the high level.
@@
Our controller is designed to function as a low-level component within this modular control hierarchy. It translates high-level total thrust and angular velocity commands into individual motor speeds with adaptation to different quadcopters. The implementation of low-level adaptation abstracts away the physical complexities of the system from the high-level planner, allowing it to focus on high-level mission tasks without concern for the intricacies of the underlying hardware. This flexibility also enables our controller to enhance non-adaptive high-level controllers, adding adaptability to disturbances and model mismatches, and thereby improving overall system performance.
II-B Learning an Adaptive Controller
An overview of our training strategy is illustrated in Figure 2. The controller learns to control randomly generated quadcopters to track diverse trajectories, which are generated with randomized motion primitives [mueller2015trajectory]. The training of the controller employs a dual strategy, combining reinforcement learning (RL) and behavior cloning (BC) from an expert model-based controller. We will provide details about each component during the training process in the following subsections.
II-B1 Reinforcement Learning
As in our previous conference work [zhang2023universal], we use RMA [kumar2021rma] for the reinforcement learning part of our training. We do not make any changes to this part from our previous conference work, but review it here for completeness.
@@ II-B1 Reinforcement Learning
Our controller consists of a base policy , an intrinsics encoder , and an adaptation module . At time , the base policy takes the current state and the ground-truth intrinsics vector to output the target motor speeds for all individual motors. The intrinsics vector is a low dimensional encoding of the environment parameters , which consist of model parameters or external disturbances that are key to adaptive control. We use the intrinsics encoder to compress to . This gives us:
@@
The current state includes the attitude matrix (), the mass-normalized thrust (), angular velocity (), commanded total thrust () and commanded angular velocity (). The environmental parameter includes mass, arm length, propeller constants (torque-to-thrust ratio and thrust-to-motorspeed-squared ratio), the diagonal entries of MMOI matrix (), body drag coefficients (), maximum motor rotation, motor effective factors (), mixer matrix (), payload mass, and external torque (), which results in an 34 dimensional vector.
@@
The low dimensionality (8) of is determined empirically through a binary search on the dimension of , aiming to optimize the learning performance of the policy. The latent representation of high-dimensional system parameters allows the base policy to adapt to variations in drone parameters, payloads, and disturbances such as external force or torque.
@@
During deployment, we do not have access to and hence we cannot directly measure in the real world. Instead, we estimate it via the adaptation module , which uses the commanded actions and the measured sensor readings from the latest steps to estimate it online during deployment as Equation 3. We can train this adaptation module in simulation using supervised learning because we have access to the ground truth intrinsics . We minimize the mean squared error loss when is estimated using sensor-action history of the vehicle tracking online generated random trajectories. The random trajectory tracking task can provide a set of rich excitation signals for the adaptation module to estimate using the sensor-action history.
@@
The estimated along with the current state is fed into our base policy to output motor speed during deployment as Equation 4. More concretely,
II-B2 Guiding search by imitating a model-based controller
We implement the base policy and the intrinsics encoder as Multi-layer perceptrons and jointly train them end-to-end in simulation. The training is done by an integration of BC and model-free RL. In our approach, the expert controller is a model-based low-level controller. The key distinction from previous work combining reinforcement and imitation learning [xie2020learning, huang2024manipulator, loquercio2023learning] is that during each training episode, the ground-truth model parameters of randomized quadcopters are used to adapt the expert controller. This ensures that our base policy learns from an expert controller that dynamically adjusts its behavior whenever the quadcopter model changes. The BC loss minimizes the mean squared error loss on actions:
Similarly to previous work, we adaptively change the relative weight between these two losses. The weight of the BC losses decays exponentially while the weight for RL increases inversely with training steps so that RL becomes dominant later in the training process. This training scheme enables rapid learning of the desired behavior from the expert controller at the beginning of training and generalization by RL in the later parts of training. The overall training framework seeks to maximize the overall reward of the policy :
This learning approach improves the training of the base policy compared to training solely with RL, as in our previous conference paper [zhang2023universal]. Figure 3 presents an ablation of the BC component of the loss. The policy trained with both BC and RL shows steady improvement in reward and episode length throughout the learning phase, maintaining an episode length close to 1 after approximately 10 million steps. In contrast, the RL-only baseline improves at a slightly slower rate, achieving reward similar to the BC-and-RL method at around 50 million steps. However, the RL-only baseline does not perform as consistently and begins to diverge after 50 million steps, losing much of its performance gains and eventually leading to crashes.
II-B3 Reward Design
@@
Output Command Smoothing Penalty:
@@
The output smoothing penalty discourages high-frequency control commands, which can overheat the quadcopter’s motors and damage the hardware system during deployment. The survival reward encourages the quadcopter to learn to fly longer until the end of the training episode. Finally, both tracking deviation penalties encourage the quadcopter to track the given high-level commands by matching its mass-normalized thrust and torque with the reference commands.
While angular velocity is one of the high-level commands to track and one of our policy’s observations during deployment, we prioritize rewarding torque tracking in our learning framework. Our experiments indicate that this design choice leads to noticeable improvements in learning speed, stability, and overall reward. Torque responds directly and immediately to applied motor speed commands, whereas angular velocity is derived through sequential integration of the commanded motor speed. Therefore, the torque tracking reward design can provide more direct reward after a corrective action, which is important for training a low-level high-bandwidth controller like ours, with such high frequency at 500Hz.
We compare the training curves of the policy trained with torque tracking and that with angular velocity tracking as shown in Figure 4, keeping all other reward design and hyperparameters the same. Specifically, we calculate the total reward per episode, excluding the torque / angular velocity tracking penalty, to evaluate the performance of the two policies. The torque tracking baseline improves throughout the training time steps, whereas the reward of the angular velocity tracking baseline grows with larger fluctuations and eventually ends at a lower reward.
II-B4 Quadcopter Randomization
@@
The training of our adaptive policy requires a wide spectrum of quadcopters, a challenge that we address through a carefully crafted randomization process. We propose a randomization method that embodies key physical principles and design constraints of quadcopters, ensuring that the generated variations are physically plausible. Quadcopters follow a general design pattern, which typically involves a symmetric structure with four rotors positioned at the corners of a square frame. The size of a quadcopter is positively correlated with its mass, moment of inertia, and other properties, such as motor power and body drag coefficient. Our method follows the pattern in randomizing the quadcopters and their respective dynamic characteristics, instead of simply varying parameters independently. In particular, we introduce a few key factors in quadcopter randomization which govern the variation of some other quadcopter body parameters.
@@
Assuming a constant density and proportional scaling in all dimensions, the mass of the quadcopter is directly proportional to its volume, which in turn scales with the cube of its arm length. Similarly, the moment of inertia, which depends on both the mass distribution and the distance from the axis of rotation, scales approximately with the fifth power of the arm length under these assumptions. The body drag coefficient, primarily influenced by the cross-sectional area the quadcopter presents to the airflow, scales with the square of the arm length.
@@
Noise. To ensure flexibility and avoid adhering too strictly to the scaling rule, we introduce a uniformly distributed noise in the range of [-20%, 20%] to all parameters after they have been scaled with the size factor .
Motor Efficiency Factor. We randomize the motor efficiency factor for each of the four rotors. For each rotor, the simulated motor speed is calculated by multiplying the intended speed by this factor. This is to simulate the motor ineffectiveness due to battery voltage drop, a partial motor failure, or simply hardware variations.
External Disturbance. At a randomly sampled time during each episode, the parameters, including mass, inertia, and the center of mass, are again randomized. This is used to mimic sudden variations in the quadcopter parameters due to a sudden disturbance caused by an off-center payload or wind.

### v2

# A Learning-based Quadcopter Controller with Extreme Adaptation

@@ I Introduction
The agile nature of quadcopters and the necessity for precise control in dynamic environments create a unique context for exploring control strategies. Model-based controllers for quadcopters generally rely on estimates of the vehicle’s properties, including inertia, motor constants, and other parameters. Notable examples include sliding mode control [1] and PID controllers [2]. Once these parameters are estimated, the controller typically requires iterative tuning through successive experiments to refine its performance. However, inaccuracies in parameter estimation can directly lead to execution errors in controller commands. Furthermore, any modification to the vehicle, such as attaching an extra payload, could lead to suboptimal performance without repeating the estimation and tuning processes. Such significant engineering effort could be eased with a universal controller that does not require specialized tuning.
In this work, we propose a learning-based low-level controller designed to control a variety of quadcopters with notable differences in mass, size, propellers, and motors. Our controller is also capable of rapidly adapting to unknown in-flight disturbances such as off-center payloads, loss of efficiency in motors, and wind.
@@ I-A Related Work
The development of fixed-parameter controllers, while foundational, is inherently limited by their lack of real-time adaptivity to model uncertainties and disturbances. Adaptive control techniques were introduced to address these unpredicted variations in a system. One of the initial contributions in this field was the model reference adaptive controller (MRAC), an extension of the well-known MIT-rule [3]. The empirical success of MRAC led to the development of control [4, 5], which offers a promising solution by estimating the differences between the nominal state transitions predicted by the reference model and those observed in practice. Such differences are compensated by allocating a control authority proportional to the disturbance, effectively driving the system back to its reference behavior.
@@
However, the performance of the classic formulation degrades when the observed transitions deviate greatly from the (usually linear) reference model. This limitation is critical in scenarios involving large variations between different quadcopters’ dynamics, as the underlying assumptions on locally linear disturbances are generally not fulfilled. Some studies have discussed practical considerations regarding its implementation. For instance, high adaptive gains in certain scenarios may introduce numerical challenges and sensitivity to unmodeled dynamics [6]. It has also been noted that, under specific conditions, the closed-loop behavior of controllers can resemble that of linear proportional-integral (PI) controllers [7, 8], which may limit the expression of its adaptive features. To address these issues, recent research has explored combining adaptive control with nonlinear online optimization techniques, such as model predictive control (MPC) [9, 10, 11]. These methods have achieved impressive results, but still require explicit and accurate knowledge of the reference model for adaptation.
Beyond , advanced nonlinear adaptive controllers have been developed to enable agile quadcopter maneuvers in the presence of uncertainty. Notable approaches include geometric adaptive tracking control on SE(3) [12] and adaptive incremental nonlinear dynamic inversion (INDI) [13]. These controllers improve performance by directly incorporating nonlinear dynamics and real-time uncertainty estimation into the control design. A nominal model prior is first estimated and then used for subsequent compensation of dynamic variations, enhancing robustness during high-precision maneuvers.
@@
Recent advances in data-driven control strategies have shown promising results for quadcopter stabilization [14, 15], or waypoint tracking flight [16, 17], as well as agile racing against human pilots [18]. Model-free reinforcement learning in [18] has demonstrated impressive adaptability to unmodeled disturbances, such as blade flapping effects. Combining data-driven methods with model-based control designs has also been proposed to leverage the guaranteed adaptivity and robustness offered by model-based control. For instance, some studies have learned policies from model-based methods like Robust-Tube MPC through imitation learning [19, 20, 21]. Another approach is augmenting the learned controller with classical adaptive control designs during deployment for fast disturbance estimation and online adaptation [22, 23]. Despite these advancements, these methods remain tailored to specific platforms. Transferring the same controller to another vehicle typically requires retraining or fine-tuning the policy, along with data collection for the new vehicle.
@@
Zero-shot adaptation across different vehicles has been demonstrated on quadrupeds [24], highlighting the versatility of learning-based methods. However, this generalized control relies on existing internal motor control loops rather than directly adapting at the motor level. In the case of quadcopters, the variation in actuators between different vehicles is particularly significant, with motor constants potentially differing by orders of magnitude. Consequently, effective adaptation across different quadcopter platforms must address motor-level differences directly. Additionally, the high-frequency nature of motor control increases the risk of crashes due to inadequate adaptation. These factors, the substantial variation in actuators and the high-frequency nature of motor control, pose significant challenges to applying previous methods of adaptive trajectory control to the problem of extreme adaptation across quadcopters.
@@ I-B Our Contributions
In this work, we present a general framework for learning low-level adaptive controllers that are effective across a wide range of quadcopters. Similar to prior research in learning-based control for aerial robots [25, 17, 26, 27, 14, 28], we train the policy entirely in simulation using reinforcement learning and deploy it directly to the real world without fine-tuning (i.e., zero-shot deployment).
While previous works typically rely on slight parameter randomizations (up to 30%) around a nominal model [17], our approach must handle parameter variations thousands of times larger. This presents a significant challenge for reinforcement learning, as such a broad range of variations can hinder optimization convergence.
To address this challenge, we build on our earlier conference work on learning-based low-level control [29] and introduce three key technical innovations: (1) a dual training strategy that combines imitation learning from specialized model-based controllers and model-free reinforcement learning. This combination effectively handles the challenges of training a low-level controller due to its high-frequency nature and the low informational density of observations. (2) A specifically designed reward to provide the low-level controller with direct feedback for quick adjustments, allowing it to perform more agile maneuvers, and (3) a designed-informed domain randomization method to ensure that the variations in quadcopter designs during training are consistent with real-world constraints. These innovations eliminate constraints on slow flight and accurate state estimation, and widens the range of vehicles our policy can fly. Our approach significantly outperforms existing baselines. In addition, it enables the controller to adapt to out-of-distribution quadcopters in simulation up to 16x wider than the training set and to disturbances for which it was not explicitly trained, such as wind.
@@ II Method
We present our methodology for learning an adaptive low-level controller for various quadcopters. A list of symbols and notations are given in Table I.
Notation (for arbitrary symbol ) | |||
@@
Commanded quantity | Quantity in reference trajectory | ||
@@
Individual motor speeds | Individual motor Pulse width modulation (PWM) commands | ||
@@ II-A Control Structure
Cascade control systems are instrumental in managing complex dynamic systems by decomposing them into a hierarchy of simpler nested subsystems. For example, in a two-layer system, the high-level component focuses on high-level tasks such as trajectory planning, while the low-level component acts as the inner control loop to execute the commands from the high level.
@@
Our controller is designed to function as a low-level component within this modular control hierarchy. It translates high-level total thrust and angular velocity commands into individual motor speeds with adaptation to different quadcopters. The implementation of low-level adaptation abstracts away the physical complexities of the system from the high-level planner, allowing the latter to focus on high-level mission tasks. This flexibility also enables our controller to enhance non-adaptive high-level controllers, adding adaptability to disturbances and model mismatches and thereby improving overall system performance.
An overview of our training strategy is illustrated in Figure 2. In simulation, the controller learns to track diverse trajectories across various quadcopter models. The training follows a dual approach, combining reinforcement learning (RL) with imitation learning (IL) from an expert model-based controller. To ensure efficient and realistic training, we employ a design-informed randomization law to sample quadcopter parameters. This approach maintains adherence to real-world design constraints while enhancing training efficiency. The generated parameters are then used to update the expert model-based controller, ensuring it adapts its behavior as the simulated model changes.
The following subsections provide details on each component of the training process.
II-B Training Framework
We adopt a learning-based framework that decouples policy learning into a control policy and a real-time estimator, commonly used in quadruped locomotion [30, 31]. The detailed training process is shown in Figure 3. We train in two phases. In the first, we train a low-level controller given access to ground-truth system parameters via RL and IL. Since we don’t have such parameters in the real world, we use a second phase to learn an adaptation module . Such a module predicts the system parameter from a sensor-action history. The module is trained in simulation using supervised learning. During deployment, we can use the base policy and the adaptation module to achieve zero-shot adaptation.
@@ II-B Training Framework
Phase 1 Training. Our controller consists of a base policy , an intrinsics encoder , and an adaptation module . At time , the base policy takes the current state and the ground-truth intrinsics vector to output the target motor speeds for all individual motors. The intrinsics vector is a low dimensional encoding of the environment parameters , which consist of model parameters or external disturbances that are key to adaptive control. We use the intrinsics encoder to compress to . This gives us:
@@
The current state includes the mass-normalized thrust (), angular velocity (), commanded total thrust () and commanded angular velocity (). The environmental parameter includes mass , arm length , propeller constants (torque-to-thrust ratio and thrust-to-motorspeed-squared ratio ), the diagonal entries of MMOI matrix (), body drag coefficients (), maximum motor rotation, motor effective factors (), mixer matrix (), payload mass, and external torque (), which results in an 34 dimensional vector. The mixer matrix is the inverse of the allocation matrix. While the allocation matrix maps rotor forces to total thrust and torques, the mixer matrix maps desired thrust and torques to individual rotor forces, making it essential for motor control to determine the commands for individual rotors. A more detailed derivation is provided, for example, in [32].
@@
The choice of 8 dimensions for the intrinsics vector is determined empirically through a binary search on the dimension of the intrinsics , aiming to optimize the learning performance of the policy. The latent representation of high-dimensional system parameters allows the base policy to adapt to variations in drone parameters, payloads, and disturbances such as external force or torque.
@@
Phase 2 Training. During deployment, we do not have access to the environmental parameters and hence we cannot directly measure the intrinsics in the real world. Instead, we estimate it via the adaptation module , which uses the commanded actions and the measured sensor readings from the latest steps to estimate it online during deployment as (3). We can train this adaptation module in simulation using supervised learning because we have access to the ground truth intrinsics . We minimize the mean squared error loss when is estimated using sensor-action history of the vehicle tracking trajectories generated with randomized motion primitives [33]. The random trajectory tracking task can provide a set of rich excitation signals for the adaptation module to estimate using the sensor-action history.
@@
The estimated along with the current state is fed into our base policy to output motor speed during deployment as (4). More concretely,
II-C Reward Design
@@
Output Smoothness Penalty:
@@
The output oscillation penalty encourages smooth inputs, discouraging high-frequency control signals. The survival reward encourages the quadcopter to learn to fly longer until the end of the training episode. Finally, both tracking deviation penalties encourage the quadcopter to track the given high-level commands by matching its mass-normalized thrust and torque with the reference commands.
II-D Guiding search by imitating a model-based controller
We implement the base policy and the intrinsics encoder as multi-layer perceptrons and jointly train them end-to-end in simulation. The training is done by an integration of IL and model-free RL. In our approach, the expert controller is a low-level proportional-derivative (PD) controller with access to the sampled vehicle’s ground truth model parameters. It receives mass-normalized thrust and angular velocity commands from the high-level controller and computes the desired torque using (6). The resulting thrust and torque are linearly mapped to individual motor forces () using the mixer matrix (), which are then converted into motor speeds (), representing the expert action.
where the square root operation is applied elementwise.
The key distinction from other work combining reinforcement and imitation learning [34, 35, 36] is that during each training episode, the ground-truth model parameters of randomized quadcopters are used to adapt the expert controller. This ensures that our base policy learns from an expert controller that dynamically adjusts its behavior whenever the quadcopter model changes. The IL loss minimizes the mean squared error loss on actions:
We adaptively change the relative weight between these two losses. The overall training framework seeks to maximize the overall reward of the policy :
where the weight of the IL losses decays exponentially while the weight for RL increases inversely with training steps so that RL becomes dominant later in the training process. This training scheme enables rapid learning of the desired behavior from the expert controller at the beginning of training and generalization by RL in the later parts of training.
II-E Quadcopter Parametric Randomization
@@
The training of our adaptive policy requires a wide spectrum of quadcopters, a challenge that we address through a carefully crafted randomization process rather than relying on uniform sampling. We propose a randomization method that embodies key physical principles and design constraints of quadcopters, ensuring that the generated variations are physically plausible. Quadcopters follow a general design pattern, which typically involves a symmetric structure with four rotors positioned at the corners of a square frame. The size of a quadcopter is positively correlated with its mass, moment of inertia, and other properties, such as motor power and body drag coefficient. Our method follows the pattern in randomizing the quadcopters and their respective dynamic characteristics, instead of simply varying parameters independently. In particular, we introduce a few key factors in quadcopter randomization which govern the variation of some other quadcopter body parameters.
@@
Assuming a constant density and proportional scaling in all dimensions, the mass of the quadcopter is proportional to its volume, which in turn scales with the cube of its arm length. Similarly, the moment of inertia, which depends on both the mass distribution and the distance from the axis of rotation, scales approximately with the fifth power of the arm length under these assumptions. The body drag coefficient, primarily influenced by the cross-sectional area the quadcopter presents to the airflow, scales with the square of the arm length.
@@
Noise. To ensure flexibility and allow for the fact that real systems will not perfectly follow these scaling rules, we introduce a uniformly distributed noise in the range of [-20%, 20%] to all parameters after they have been scaled with the size factor .
Motor Effectiveness Factor. We randomize the motor effectiveness factor for each of the four rotors so that the motor will produce a force different than expected. For each rotor, the simulated motor speed is calculated by multiplying the intended speed by this factor. This is to simulate the motor ineffectiveness due to battery voltage drop, a damaged propeller, or simply hardware variations.
External Disturbance. At a randomly sampled time during each episode, the parameters, including mass, inertia, and the center of mass, are again randomized. This is used to mimic sudden variations in the quadcopter parameters due to a sudden disturbance caused by an off-center payload.